OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 3x – y = 2 6x – 2y = 4

OpenStudy (anonymous):

do the same thing as the one i just answered

OpenStudy (anonymous):

this site is for help not for people to do everything for you

OpenStudy (anonymous):

If Yu Dnt Like It Den Bye !

OpenStudy (anonymous):

If you are not interested in learning how to do it and doing it yourself, feel free to plug in your system to wolfram alpha and let it give you the answer.

hero (hero):

Oh my!

OpenStudy (chaise):

3x – y = 2 6x – 2y = 4 3x-2 = y 6x - 2(3x-2)=4 6x - 6x+4 = 4 4 = 4 I'm confused?

hero (hero):

Infinite Solutions

OpenStudy (anonymous):

NO SOLUTIONS!

OpenStudy (anonymous):

NOOO all real numbers

hero (hero):

Spykez, it would only be no solutions if it wasn't true.

hero (hero):

4 = 4 is true

OpenStudy (chaise):

Cool. Just needed verification :)

hero (hero):

It's the same equation

hero (hero):

Solve both equations for y and you'll see

OpenStudy (anonymous):

So Its No Solution

OpenStudy (anonymous):

NO the answer is all real numbers satisfy these equations

hero (hero):

Just because your computer says no solutions, doesn't make it correct

hero (hero):

It only thinks there's no solutions because x was eliminated when you tried to solve for x.

hero (hero):

Most programs are set up where if x is eliminated when attempting to solve for x using the substitution method, then there is no solution. But in this case, there is clearly infinite solutions.